OpenStudy (anonymous):

Is this system consistent or inconsistent? 2x + y = 5 -4x + 6y = 12

OpenStudy (tkhunny):

Have you attempted to find solutions? It almost always helps to simplify where possible. That second equation can be prettied up a bit. 2x + y = 5 -2x + 3y = 6 Now, you can just add the two equations and see what gives.

OpenStudy (anonymous):

wait did you divide the second equation by two?

OpenStudy (tkhunny):

Each term had a factor of 2. It seemed a good idea if my goal was to simplify my life.

OpenStudy (anonymous):

okay let me solve

OpenStudy (anonymous):

I still don't get it my 2x cancelled but I don't think im suppose to get a remainder

OpenStudy (unklerhaukus):

fractions are fine

OpenStudy (anonymous):

okay good! so once I get my fraction then what do I do?

OpenStudy (unklerhaukus):

you have found y?, well then, sub it into one of those equations at the top and find x

OpenStudy (anonymous):

y is 2 3/4. so how do I know if is consistent or dependent?

OpenStudy (unklerhaukus):

yes that is y , what is x? if there is at least one solution (x,y) then the system consistent if there are no possible solutions the system is inconsistent

OpenStudy (anonymous):

x is 1.125 or 1.12 1/2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

so these answers would be a solution?

OpenStudy (tkhunny):

Consistent: ANY solution. Inconsistent: NO Solution.

OpenStudy (unklerhaukus):

yes x is 1.125

OpenStudy (anonymous):

so its consistent because I found an answer?

OpenStudy (unklerhaukus):

that's it

OpenStudy (anonymous):

okay thank you :)

OpenStudy (unklerhaukus):

if you had a system like x + y = 1 x + y = 0 taking the difference 0 = 1 ??? this system is inconsistent because no matter what values x and y are , both equations cannot be satisfied simultaneously

OpenStudy (anonymous):

I understand it now thank you! how do find out if a system is dependent or not? like 2x + y = 4 and -6x - 3y = -12

OpenStudy (unklerhaukus):

try to find a solution like last time, what happens

OpenStudy (anonymous):

both the variable cancel ou

OpenStudy (anonymous):

out

OpenStudy (anonymous):

and we are

OpenStudy (anonymous):

eft

OpenStudy (anonymous):

left

OpenStudy (anonymous):

left with 0 = 0

OpenStudy (unklerhaukus):

Yes, that is right

OpenStudy (unklerhaukus):

both equations are the equation to the same straight line

OpenStudy (anonymous):

so there are infinitely many solutions so its dependent

OpenStudy (anonymous):

can you do one more?

OpenStudy (anonymous):

-x + 4y = 24 -2x + 8y = 8

OpenStudy (unklerhaukus):

-x + 4y = 24 -2x + 8y = 8

OpenStudy (unklerhaukus):

you might start by doubling the first equation

OpenStudy (anonymous):

okay ill get right on that

OpenStudy (anonymous):

answer is 0 = 56 not dependent

OpenStudy (unklerhaukus):

0 = 56 inconsistent

OpenStudy (anonymous):

confused

OpenStudy (unklerhaukus):

it doesn't make sense to say wether a system is dependent or not if is isn't even consistent

OpenStudy (unklerhaukus):

I think we should back to the last one (the dependent system), because you got the right answer but i think the reasoning was incorrect

OpenStudy (unklerhaukus):

2x + y = 4 -6x - 3y = -12 this system reduced to y = -2x+4 0 = 0 yes there are infinite solutions , (but some systems with infinite solutions are dependent) the system is dependant because y is a function of x y(x)=-2x+4 ___ if we had a system x=x y=0 this system also has infinite solutions, but this is an independent system this system is independent because y is not a function of x

OpenStudy (anonymous):

oh okay thank you I get it